# Advances in Feed-Forward 3D Reconstruction and View Synthesis: A Survey

### v3

@@ 4.1 Datasets
Datasets are the core of feed-forward 3D reconstruction and view synthesis. To give an overall picture of the datasets, we tabulate detailed scene and annotation types in popular datasets in Table I. The scene types are divided into objects, indoor and outdoor scenes. And we also indicate synthetic datasets (e.g., Objaverse [46]), where MegaSynth [255] and Zeroverse [47] are procedurally synthesized datasets, real-world datasets (e.g., ACID [256]), static datasets (e.g., ARKitScenes [257]) and dynamic datasets (e.g., KITTI360 [258]). Notably, several datasets, for example TartanAir [259], include both static and dynamic scenes.
TABLE I: Summarization of popular datasets for feed-forward 3D reconstruction and view synthesis.
DTU [260] 124 Objects Real
Pix3D [261] 395 Objects Real
GSO [262] 1,030 Objects Real
OmniObject3D [263] 6,000 Objects Synthetic
CO3D [27] 18,619 Objects Real
WildRGBD [264] 23,049 Objects Real
ShapeNet [265] 51,300 Objects Synthetic
MVImgNet [266] 219,188 Objects Real
Zeroverse [47] 400K Objects Synthetic
Objaverse [46] 818K Objects Synthetic
Objaverse-XL [267] 10.2M Objects Synthetic
7Scenes [268] 7 Indoor Scenes Real
Replica [269] 18 Indoor Scenes Real
TUM RGBD [270] 39 Indoor Scenes Real
Matterport3D [271] 90 Indoor Scenes Real
HyperSim [272] 461 Indoor Scenes Synthetic
Dynamic Replica [273] 524 Indoor Scenes Synthetic
ScanNet++ [274] 1,006 Indoor Scenes Real
ScanNet [275] 1,513 Indoor Scenes Real
ARKitScenes [257] 1,661 Indoor Scenes Real
MegaSynth [255] 700K Indoor Scenes Synthetic
Virtual KITTI2 [276] 5 Outdoor Scenes Synthetic
KITTI360 [258] 11 Outdoor Scenes Real
Spring [277] 47 Outdoor Scenes Synthetic
MegaDepth [278] 196 Outdoor Scenes Real
ACID [256] 13,047 Outdoor Scenes Real
MipNeRF360 [279] 9 Indoor and Outdoor Scenes Real
Tanks&Temples [280] 21 Indoor and Outdoor Scenes Real
ETH3D [281] 25 Indoor and Outdoor Scenes Real
PointOdyssey [282] 159 Indoor and Outdoor Scenes Synthetic
TartanAir [259] 1,037 Indoor and Outdoor Scenes Synthetic
DL3DV-10K [283] 10,510 Indoor and Outdoor Scenes Real
RealEstate10K [16] 74,766 Indoor and Outdoor Scenes Real
BlendedMVS [46] 113 Objects, Indoor and Outdoor Scenes Synthetic
@@ 4.2 Evaluation Metrics
Several metrics have been widely adopted for faithful evaluations in various feed-forward 3D reconstruction and view synthesis tasks. For novel view synthesis evaluation, PSNR (Peak Signal-to-Noise Ratio), SSIM (Structural Similarity Index) [284], and LPIPS (Learned Perceptual Image Patch Similarity) [285] are commonly used to evaluate image quality from different perspectives.
@@
For dynamic point tracking, OA (Occlusion Accuracy), $\sigma^{vis}_{avg}$, and AJ (Average Jaccard) [286] are used together. OA measures the binary accuracy of the occlusion predictions; $\sigma^{x}_{avg}$ measures the fraction of points that are accurately tracked within a certain pixel threshold; Average Jaccard considers both occlusion and prediction accuracy.
@@ 5.1 Limited Modality in Datasets
Most existing 3D reconstruction and view synthesis datasets have a limited coverage of data modalities. Specifically, many widely-used benchmarks, such as RealEstate10K [16] and MipNeRF360 [279], comprise RGB images only without including essential complementary signals like depth, LiDAR, or semantic annotations. Even large-scale collections like Objaverse-XL [267] (10.2M objects) focus primarily on synthetic mesh data, lacking the real-world data modalities needed to train robust models. Many studies address this imbalance issue by merging multiple datasets of different modalities, but this inevitably introduces domain shifts and annotation inconsistencies. The modality limitation is particularly acute in the area of dynamic scene understanding. While several datasets provide dynamic sequences, those with comprehensive multi-modal annotations (e.g., synchronized RGB, depth, optical flow, and 3D motion) remain significantly fewer than their static counterparts. Most dynamic datasets prioritize either camera motion or object movement, but rarely capture both simultaneously with full sensor suites. This scarcity of richly annotated dynamic data severely constrains the development of models capable of handling real-world scenarios that often involve both camera motion and object motion.
A fundamental challenge emerges: how to create scalable, modality-rich datasets that combine the diversity of synthetic collections like Objaverse [46] with the multi-sensor completeness of real-world benchmarks such as ScanNet++ [274]. Current approaches address this issue by patching together incompatible data sources, which ultimately limits progress toward generalizable 3D understanding. The field is facing an urgent need of comprehensive resources that provide aligned multi-modal signals, including RGB, depth and semantics, all collected under a unified protocol for mitigating the data modality limitation.
@@ 5.2 Reconstruction Accuracy
Feed-forward 3D reconstruction models have made notable progress in recent years. However, their reconstruction accuracy, particularly in terms of depth map precision, is still inferior to traditional multiview stereo (MVS) methods [287, 15, 288] that explicitly utilize camera parameters for all input frames. Specifically, MVS approaches typically leverage known camera parameters and hypothesized depth sets to construct cost volumes, subsequently processed to predict accurate depth or disparity maps. An intriguing hypothesis is that feed-forward 3D reconstruction models might spontaneously learn an approximation of such cost volumes. Modern feed-forward reconstruction models [2, 59] mostly employ self-attention layers, theoretically enabling them to approximate or even exceed the representational capacity of traditional cost volumes. With sufficient high-quality training data, these feed-forward models have the potential to match and even surpass the accuracy of MVS-based methods. Moreover, incorporating explicit camera parameters or additional priors into the feature backbone, such as through Diffusion Transformers [289], offers another promising avenue to enhance reconstruction accuracy. Consequently, we anticipate that feed-forward models will continue to evolve, eventually much outperforming traditional MVS methods and achieving sensor-level accuracy, comparable to technologies like LiDAR or high-precision scanning systems.
@@ 5.4 Long Context Input
Existing methods for 3D geometry reasoning and novel view synthesis often rely on full attention mechanisms, which lead to a cubic increase in token count and computational cost. For example, inferring from 50 images with VGGT [2] requires approximately 21 GB of GPU memory, while scaling to 150 images - even with advanced techniques like FlashAttention2 [290] - demands around 43 GB. Training on more than 32 views remains infeasible even on the most powerful GPUs. A promising alternative is the use of recurrent mechanisms, such as Cut3R [3], which incrementally integrate new views while maintaining a memory state. Although this approach keeps inference memory usage consistently low (i.e., around 8 GB in practice), it suffers from forgetting previously seen information, leading to significant performance degradation as the number of input views increases. Efficiently reasoning over hundreds or even thousands of views while keeping memory and computation costs manageable remains an open and pressing challenge.
@@ 6.1 Applications
3D reconstruction models have a wide range of applications with positive societal impacts. To name a few, they have the potential to transform the film and gaming industries with more realistic visual effects and production speed by using reconstructed or generated 3D assets. They are also valuable in the development of smart cities, where they can be used to create ”digital twins” of critical infrastructure for simulation and maintenance planning. Additionally, 3D reconstruction can help cultural heritage preservation, as it allows ancient artifacts and statues to be digitally preserved before they deteriorate.
@@ 6.2 Misuse
The widespread availability of 3D reconstruction models could introduce various misuses. One typical concern is related to privacy. For example, private property could be reconstructed without the owner’s permission simply by taking a few pictures. To address this issue, new regulations should be established as 3D reconstruction technologies become increasingly accessible. In addition, the generative capabilities of feed-forward 3D reconstruction models can be misused to create false evidence, such as fabricated crime scenes. To handle such misuse, advanced detection models should be developed that can distinguish between generated and real content. People can also develop techniques to add ”invisible watermarks” on generated outputs, allowing simple decoding to verify if content is artificially created.

### v4

@@ 4.1 Datasets
Datasets are the core of feed-forward 3D reconstruction and view synthesis. To give an overall picture of the datasets, we tabulate detailed scene and annotation types in popular datasets in Table 1. The scene types are divided into objects, indoor and outdoor scenes. And we also indicate synthetic datasets (e.g., Objaverse [DSS23]), where MegaSynth [JXX25] and Zeroverse [XBS24] are procedurally synthesized datasets, real-world datasets (e.g., ACID [LTJ21]), static datasets (e.g., ARKitScenes [BCD21]) and dynamic datasets (e.g., KITTI360 [LXG22]). Notably, several datasets, for example TartanAir [WZW20], include both static and dynamic scenes.
@@ 4.2 Evaluation Metrics
Several metrics have been widely adopted for faithful evaluations in various feed-forward 3D reconstruction and view synthesis tasks. For novel view synthesis evaluation, PSNR (Peak Signal-to-Noise Ratio), SSIM (Structural Similarity Index) [WBSS04], and LPIPS (Learned Perceptual Image Patch Similarity) [ZIE18] are commonly used to evaluate image quality from different perspectives.
@@
For dynamic point tracking, OA (Occlusion Accuracy), $\sigma^{vis}_{avg}$, and AJ (Average Jaccard) [DGM22] are used together. OA measures the binary accuracy of the occlusion predictions; $\sigma^{x}_{avg}$ measures the fraction of points that are accurately tracked within a certain pixel threshold; Average Jaccard considers both occlusion and prediction accuracy.
4.3 Evaluation Results
Tables 2-5 summarize quantitative results across major benchmarks, covering camera pose estimation, point map reconstruction, video depth estimation, and novel view synthesis. These comparisons demonstrate the rapid evolution of feed-forward 3D reconstruction models in terms of accuracy and scalability.
Camera Pose Estimation. As shown in Table 2, recent transformer-based methods (e.g., VGGT [WCK25] and $\pi^{3}$ [WZZ25]) achieve state-of-the-art performance on Sintel, TUM-dynamics, ScanNet, RealEstate10K, and Co3Dv2. Compared to early baselines such as DUSt3R and Spann3R, these large-scale models reduce both translation and rotation errors by more than 50% while maintaining high accuracy under wide-baseline conditions. Notably, PAGE-4D [ZWC25] improves pose estimation in dynamic scenes by disentangling static and dynamic cues, demonstrating strong robustness for real-world motion sequences.
Point Map Estimation. Table 3 reports evaluations on 7-Scenes and NRGBD datasets. VGGT achieves the best overall results, attaining mean accuracy below 0.09 and the highest normal consistency (0.89), confirming the effectiveness of large transformer architectures in maintaining geometric fidelity. Feed-forward stream-based approaches such as CUT3R and STream3R surpass optimization-based pipelines (DUSt3R-GA, MASt3R-GA) in both accuracy and normal consistency, without causing a large burden with the continuous updating mechanisms.
Video Depth Estimation. In Table 4, models are evaluated on Sintel, Bonn, and KITTI. Under both scale-aligned and scale-shift aligned protocols, $\pi^{3}$ again delivers the most accurate depth with the lowest absolute relative error (0.21–0.23) and highest inlier ratio ($\delta<1.25$ up to 0.99). VGGT and CUT3R follow closely, while MVDUSt3R+ demonstrates stable performance across varying numbers of input views. These results suggest that feed-forward models can generalize to complex video geometry without per-sequence optimization.
Novel View Synthesis. For single-image novel view synthesis (Table 5), generative diffusion-based approaches have advanced image quality on Tanks-and-Temples, RealEstate10K, and DL3DV datasets. PE-Fields achieves the highest PSNR (22.1–22.2) and lowest LPIPS (0.15–0.17), surpassing previous state-of-the-art methods such as SEVA and GEN3C. Recent large multiview diffusion models (e.g., See3D and MVGenMaster) show consistent improvement in perceptual realism and structural similarity, marking a clear trend toward unified feed-forward novel view synthesis and image generation.
@@ 5.1 Limited Modality in Datasets
Most existing 3D reconstruction and view synthesis datasets have a limited coverage of data modalities. Specifically, many widely-used benchmarks, such as RealEstate10K [ZTF18] and MipNeRF360 [BMV22], comprise RGB images only without including essential complementary signals like depth, LiDAR, or semantic annotations. Even large-scale collections like Objaverse-XL [DLW23] (10.2M objects) focus primarily on synthetic mesh data, lacking the real-world data modalities needed to train robust models. Many studies address this imbalance issue by merging multiple datasets of different modalities, but this inevitably introduces domain shifts and annotation inconsistencies. The modality limitation is particularly acute in the area of dynamic scene understanding. While several datasets provide dynamic sequences, those with comprehensive multi-modal annotations (e.g., synchronized RGB, depth, optical flow, and 3D motion) remain significantly fewer than their static counterparts. Most dynamic datasets prioritize either camera motion or object movement, but rarely capture both simultaneously with full sensor suites. This scarcity of richly annotated dynamic data severely constrains the development of models capable of handling real-world scenarios that often involve both camera motion and object motion.
A fundamental challenge emerges: how to create scalable, modality-rich datasets that combine the diversity of synthetic collections like Objaverse [DSS23] with the multi-sensor completeness of real-world benchmarks such as ScanNet++ [YLND23]. Current approaches address this issue by patching together incompatible data sources, which ultimately limits progress toward generalizable 3D understanding. The field is facing an urgent need of comprehensive resources that provide aligned multi-modal signals, including RGB, depth, and semantics, all collected under a unified protocol for mitigating the data modality limitation.
@@ 5.2 Reconstruction Accuracy
Feed-forward 3D reconstruction models have made notable progress in recent years. However, their reconstruction accuracy, particularly in terms of depth map precision, is still inferior to traditional multiview stereo (MVS) methods [GCS06, YLL18, ZPHW23] that explicitly utilize camera parameters for all input frames. Specifically, MVS approaches typically leverage known camera parameters and hypothesized depth sets to construct cost volumes, subsequently processed to predict accurate depth or disparity maps. An intriguing hypothesis is that feed-forward 3D reconstruction models might spontaneously learn an approximation of such cost volumes. Modern feed-forward reconstruction models [WCK25, YSL25] mostly employ self-attention layers, theoretically enabling them to approximate or even exceed the representational capacity of traditional cost volumes. With sufficient high-quality training data, these feed-forward models have the potential to match and even surpass the accuracy of MVS-based methods. Moreover, incorporating explicit camera parameters or additional priors into the feature backbone, such as through Diffusion Transformers [PX23], offers another promising avenue to improve the reconstruction accuracy. Consequently, we anticipate that feed-forward models will continue to evolve, eventually much outperforming traditional MVS methods and achieving sensor-level accuracy, comparable to technologies like LiDAR or high-precision scanning systems.
@@ 5.4 Long Context Input
Existing methods for 3D geometry reasoning and novel view synthesis often rely on full attention mechanisms, which lead to a cubic increase in token count and computational cost. For example, inferring from 50 images with VGGT [WCK25] requires approximately 21 GB of GPU memory, while scaling to 150 images - even with advanced techniques like FlashAttention2 [Dao24] - demands around 43 GB. Training on more than 32 views remains infeasible even on the most powerful GPUs. A promising alternative is the use of recurrent mechanisms, such as Cut3R [WZH25], which incrementally integrate new views while maintaining a memory state. Although this approach keeps inference memory usage consistently low (i.e., around 8 GB in practice), it suffers from forgetting previously seen information, leading to significant performance degradation as the number of input views increases. Efficiently reasoning over hundreds or even thousands of views while keeping memory and computation costs manageable remains an open and pressing challenge.
@@ 6.1 Applications
3D reconstruction models have a wide range of applications with positive societal impacts. To name a few, they have the potential to transform the film and gaming industries with more realistic visual effects and production speed by using reconstructed or generated 3D assets. They are also valuable in the development of smart cities, where they can be used to create "digital twins" of critical infrastructure for simulation and maintenance planning. Additionally, 3D reconstruction can help cultural heritage preservation, as it allows ancient artifacts and statues to be digitally preserved before they deteriorate.
@@ 6.2 Misuse
The widespread availability of 3D reconstruction models could introduce various misuses. One typical concern is related to privacy. For example, private property could be reconstructed without the owner’s permission simply by taking a few pictures. To address this issue, new regulations should be established as 3D reconstruction technologies become increasingly accessible. In addition, the generative capabilities of feed-forward 3D reconstruction models can be misused to create false evidence, such as fabricated crime scenes. To handle such misuse, advanced detection models should be developed that can distinguish between generated and real content. People can also develop techniques to add "invisible watermarks" on generated outputs, allowing simple decoding to verify if content is artificially created.